# Nanoscale magnetometry of a synthetic three-dimensional spin texture

### v1

@@ 2.3 Internal structure of the ferromagnetic stripe
To avoid having strong stray field components producing spin-mixing and be able to perform quantitative measurements, we employed a (111)-oriented NV probe. With this diamond probe we determined (see section 2 in SI). Therefore, we can assume a collinear alignment between the NV-axis and the sample magnetization, as illustrated in Figure 3(a). An NV-PL quenching map recorded without applying an external field is shown in Figure 3(b). The PL signal is quenched only along the FM stripe between AF domains. However, in contrast to the case of (100)-NV tip, the local ODMR measurements in Figure 3(c) show symmetric Zeeman splitting with respect to , on both the AF domain and the FM stripe (see Section 7 in SI for additional measurements). Based on these ODMR results, we can rule out a PL-quenching induce by spin-mixing. Thus, we attribute the NV-PL quenching to magnetic noise Finco2021-ne. From the spectra depicted in Figure 3(c), we can determine a magnetic field projected on the NV-axis of 0.6 mT and 4.0 mT on both the AF domain and the FM stripe, respectively.
To gain quantitative insight into the FM texture of the stripes observed between the AF domains in NV-PL quenching maps and MFM images, we first performed NV-SPM measurements in the iso-B imaging mode by acquiring NV-PL maps with a given MW frequency to resolve iso-magnetic field contoursRondin_2014 (see section 8 in SI). Dual iso-B images are obtained by subtracting the PL images at two fixed frequencies. In this case, we observe contour maps with two different values of the sample stray field. The dual iso-B image in Figure 3(d) reveals contours of FM domains in the nanometer range with a stray field mT and mT. The orientation of the field changes periodically, in analogy with MFM. Zero-field regions are resolved in the dual iso-B image in Figure 3(e). We observe an area of hundreds of nm2 with mT at the AF domain on the right. Additionally, there is a zero-field line between domains. Since the NV-axis is essentially perpendicular to the sample surface, this zero-field line indicates that there is no field projection along the NV-axis. Therefore, it can be interpreted as the position at which the OOP component of the field is zero, which could define the position of the DW inside the FM stripe. Tetienne2014, Tetienne2015-rc, McLaughlin2023. FM contours with mT are observed along one side of the zero-field line, while domains with opposite orientation of the field components are periodically separated by the zero-field line as shown in the image in Figure 3(f) for FM domains with mT.
@@ 2.4 Quantitative vector-field maps of a 3D spin texture
Furthermore, we record ODMR maps with a (111)-NV probe as shown in Figure 4(a). The field component projected along the NV-axis is spatially resolved over the sample surface. Given this crystallographic orientation of the tip, we are probing the = - component of the stray magnetic field. 2D maps of the in-plane (IP) stray field components can be reconstructed from the experimental map in Figure 4(a) by using the reverse propagation method in Fourier space Lima2009, dovzhenko2018 (see section 9 in SI). The obtained and stray field maps are depicted in Figure 4(b) and (c), respectively. The maximum value for the static stray field is 8.0 mT at the boundary of the AF domains. The intensity of the stray field progressively decreases when moving far away from the boundary by hundreds of nm. Both IP and OOP components of the stray field are in the range from hundreds of T to 1mT on the AF domains. Even if we expect in a multilayered SAF structure with PMA a full compensation of the magnetization on the AF domains, there will be a small field sensed by the NV above the sample surface due to the different distances between the NV center and each of the FM blocks Finco2021-ne, Kiselev2008. Additionally, from the images we observe a granular contrast, indicating magnetization variations (see section 10 in SI for a close-up ODMR map on the AF domains). These inhomogeneities or granularity are expected in sputtered multilayered samples with non-uniform thin film thicknesses or atomically flat surface that can induce local variation and non-compensation of the magnetization.
@@
Profiles taken across the region between the AF domains are plotted in Figure 4(g). For simplicity, we define the DW as the region where the OOP field component () is zero Tetienne2015-rc, McLaughlin2023, Tetienne2014 (centred at distance=0 in the profiles). Since it is the region where there is a sign reversal of the stray field (as shown in the dual iso-B results). Note that it strictly corresponds to the actual DW center line (, -magnetization) only for DWs between two similar domains (AF-AF or FM-FM). Between a FM and an AF domain, however, we only consider the DW on the top FM block. Due to the extended FM stray field, the line is shifted towards the AF domain compared to the line. At 75 nm height above the surface, this shift amounts to several tens of nanometers according to our stray field simulation. From the profiles we can also see that above the DW, the IP field components and have a local minimum and a maximum, respectively. In contrast to what has been observed in ultra-thin film Tetienne2015-rc, McLaughlin2023, the profiles of the stray field intensity are not symmetric around the DW. We attribute this to the fact that we are imaging the DW at the top FM block with a collective contribution of the full multilayered structure to the total stray field, in agreement with reported MFM profiles.Hellwig2003-prl, HELLWIG2007 The DW-type can in principle be inferred by fitting the field profiles using an analytical expression of the stray field above the DW under the approximation of a thin sample Tetienne2015-rc, i.e. a sample with magnetic thickness smaller than the sensor-sample distance. This approximation is not valid for the multilayered SAF investigated here since its magnetic thickness is comparable to the NV-sample distance. Previous micromagnetic simulations of multilayered structures demonstrated that FM domains are separated by Néel-type DWs at the top and bottom FM blocks with Bloch-type DWs in the middle of the structure Salikhov2025.
@@
Our vector-field maps are in good agreement with micromagnetic simulations. Figures 4(d)-(f) show images obtained from micromagnetic simulation of the stray field components (see methods and section 11 in SI) at a distance of 75 nm from the sample surface, in the range of the estimated . From the simulations we observed stray magnetic field of approximately 12 mT. The absolute values of the stray field are higher than the experimental values, suggesting that the magnetic parameters used to initially define the simulated structure need to be optimized as well as the determination of the NV-sample distance. However, the simulated and experimental stray-field maps are qualitatively in good agreement with the stray field values in the same order of magnitude, meaning that the remanent relaxed state achieved in the simulation can describe the magnetic texture of the multilayered SAF. Particularly, the granularity inducing a non-compensated magnetization on the AF domains is also reproduced in the simulated images. The FM domains along the stripe between the two AF domains are also well reproduced by the micromagnetic simulation. To visualize the domain structure in a 2D image, Figure 4(h) displays a transparent overlap of the simulated OOP magnetization of the top and second FM blocks (see section 11 in SI). The FM domains around the DWs are formed due to an alternating left and right lateral shift between DWs in adjacent FM blocks, leading to a vertical alignment of the magnetic moments along the stacks and giving rise to a 3D spin texture as illustrated in the inset. This lateral shift has been previously reported as driven by the dipolar interaction against the AF-exchange interaction between the FM blocks with a periodic reversal of the FM cores to minimize the magnetostatic energy produced around the DWs, and with a constant DW-shift of nm along the FM stripe between AF domains for a multilayered structure with and . Hellwig2003-prl In another study, Kiselev et al. Kiselev2010 describe, within their micromagnetic model, that the DW-shift configuration cannot be stabilized only by the interplay between the magnetostatic and the AF-exchange interaction. The stability depends also on DW pinning (grain boundaries, defects or inhomogeneities) that give rise to nucleation sites for one-dimensional FM stripes with fixed width (DW-shift). Remarkably, in the ODMR vector field maps, the periodic modulation of the stray field as a consequence of the DW-shift that reduce the magnetostatic energy is quantitatively and spatially resolved in both the experimental and the simulated IP field maps in Figure 4 (see black arrows). Additionally, the observed granularity or inhomogeneity can favor the stabilization of DW shifts through the DW pinning. We found that the lateral DW-shift is non-uniform along the FM stripes, it varies in the range nm (see section 11 in SI), generating FM domain cores with different nanometer sizes and wave-shaped DWs.
@@ 2.5 Sensing different components of the magnetic noise
Finally, in order to understand the spin dynamics of the magnetic texture in the multilayered SAF, we performed relaxometry measurements to detect spin noise. is very sensitive to magnetic fluctuations perpendicular to the NV-axis and with frequencies near the NV resonance frequency Degen2017. To investigate spin noise, we performed all-optical measurements Levine-2019 using the pulse sequences in Figure 5(a). We use a green laser pulse for initialization of the NV-spin states in , wait for some time without laser illumination, and then a second green laser pulse is sent to read out the final spin state. Figure 5(b) shows the measurements on the (111)-oriented diamond tip. We first measure before approaching the sample surface, obtaining s. This value is expected for diamond tips with shallow NV (nominal implantation depth of 10 nm), where the relaxation time is mostly affected by paramagnetic impurities on the diamond surface Rosskopf2014. After the sample approach and positioning the tip on the AF domain and then on the FM stripe, we observe a drop in by almost one and two orders of magnitude, respectively. We obtain s when the tip is in the AF domain and even shorter s on the FM stripe. Recent calculations of the spin-wave spectrum in multilayered SAFs Jimenez-Bustamante2025-ub predict GHz optical and acoustic spin-wave modes, with the dispersion relation and number of modes dependent on the thickness and number of FM blocks. Additionally, our results are in agreement with the results reported by Finco et al Finco2021-ne on the study of magnetic noise on AF DWs. Thus, we attribute these experimental observations to the presence of magnetic noise in the GHz range, most likely due to thermal magnons with frequencies near the NV spin transition frequency. is shorter on the FM stripes than on the AF domains, probably due to the confinement of thermal magnonic modes with a stronger noise or frequency that matches the electron spin resonance frequency of the NV at 2.87 GHz.
@@
Figure 5(c) shows a comparison of on NV diamond scanning probes with different crystallographic orientations (see section 12 in SI). The change in is essentially the same for the three diamond probes, indicating the existence of magnetic noise with components perpendicular to the respective NV-axis. These measurements suggest that the PL quenching observed on the FM stripes between AF domains during the acquisition of NV-PL maps could have contributions from both magnetic noise and spin-mixing. For the measurement with the (100)-NV probe, the dominant contribution to the NV-PL quenching comes probably from spin-mixing, as demonstrated by local ODMR measurements. A similar effect is expected for the (110)-NV tip. For the (111)-NV tip, the IP components resolved in the ODMR maps could in principle produce spin mixing. However, there is no evidence of spin-mixing from local ODMR measurements. Hence, the magnetic noise is the dominant contribution that produces quenching in the NV-PL images. Under this condition, it should be possible to observe an improvement in the PL contrast in the NV-PL images when the laser power is decreased below saturation, as observed on samples with AF DW Finco2021-ne. However, NV-PL quenching images recorded with different laser powers (section 13 in SI) show no improvement in PL contrast, indicating that the field of the FM core along the DW could potentially play a role in the relation between the spin-polarization rate given by the laser power and the spin relaxation rate defined by .
@@ 3 Conclusion
Quantitative vector-field maps combined with micromagnetic simulations allow us to explore interlayer coupling effects in the multilayered structure through direct magnetic imaging of DW shifts and the periodic modulation of the IP stray field components produced by the formation of nanometric FM cores around wiggled DWs. These results provide validation on an experimental and micromagnetic simulation level and further insights into existing models about the observed 3D spin texture. Additionally, local relaxometry measurements reveal the presence of GHz-range magnetic noise on both the AF domains and the DWs with FM cores.
@@ 4.1 Sample preparation and characterization
The multilayered SAF was prepared at room temperature by dc magnetron sputtering deposition (ATC 2200 series sputter tool from AJA International, Inc.) with a base pressure of Pa, using Ar sputter gas with 0.4 Pa pressure. Si substrates with a thermal 100 nm SiO2 layer were used, on which a 1.5 nm thick Ta adhesion layerEhrler2025, 20 nm Pt seed layer for promoting (111) textureZeper1991, and 3 nm Pt cap against oxidation were deposited.
@@ 4.2 NV-SPM measurements
Qualitative measurements (NV-PL maps and ) were mostly performed in a new home-built setup, which combines tuning-fork AFM (attocube tip and sample positioner and scanner) with confocal optical microscopy. NV centers in the diamond probes are optically pumped with a 515 nm laser (Toptica, iBEAM-SMART-515-S) and readout by detecting the NV-PL signal after passing through a long-pass 650 nm filter (Thorlabs, FELH0650). The signal is collected by using dual single photon counting modules (Excelitas, SPCM-AQRH-14) mounted in a Hanbury Brown and Twiss geometry for photon autocorrelation measurements. The excitation power of the green laser is 130-300 W, measured after the objective lens (Olympus, MPLFLN100x) that focus (collect) the light on (from) the tip. We use Qzabre quantum scanning tips with sensitivity Q6-Q7 and nominal NV implantation depth of 10 nm. Quantitative measurements (ODMR and dual-iso B) were carried out in a commercial Quantum Scanning NV Microscope (QSM, Qzabre AG) equipped with an electromagnet to apply magnetic field in arbitrary directions. All the NV-SPM images were acquired using the non-contact amplitude modulation mode. The NV-sample distance, as well as, the orientation of the NV-axis in the laboratory frame are determined using the method described in reference Hingant_2015, on a 1 m-wide Ta/Pt/Co(0.6nm)/Pt wire with perpendicular magnetic anisotropy as a calibration sample (see SI for more details).

### v2

@@ 2.3 Internal structure of the ferromagnetic stripe
To avoid having strong stray field components producing spin-mixing and be able to perform quantitative measurements, we employed a (111)-oriented NV probe. With this diamond probe we determined (see section 2 in SI). Therefore, we can assume a collinear alignment between the NV-axis and the sample magnetization, as illustrated in Figure 3(a). An NV-PL quenching map recorded without applying an external field is shown in Figure 3(b). The PL signal is quenched only along the FM stripe between AF domains. However, in contrast to the case of (100)-NV tip, the local ODMR measurements in Figure 3(c) show symmetric Zeeman splitting with respect to , on both the AF domain and the FM stripe (see Section 7 in SI for additional measurements). Based on these ODMR results, we can rule out a PL-quenching induce by spin-mixing. Thus, we attribute the NV-PL quenching to magnetic noise 12. From the spectra depicted in Figure 3(c), we can determine a magnetic field projected on the NV-axis of 0.6 mT and 4.0 mT on both the AF domain and the FM stripe, respectively.
To gain quantitative insight into the FM texture of the stripes observed between the AF domains in NV-PL quenching maps and MFM images, we first performed NV-SPM measurements in the iso-B imaging mode by acquiring NV-PL maps with a given MW frequency to resolve iso-magnetic field contours39 (see section 8 in SI). Dual iso-B images are obtained by subtracting the PL images at two fixed frequencies. In this case, we observe contour maps with two different values of the sample stray field. The dual iso-B image in Figure 3(d) reveals contours of FM domains in the nanometer range with a stray field mT and mT. The orientation of the field changes periodically, in analogy with MFM. Zero-field regions are resolved in the dual iso-B image in Figure 3(e). We observe an area of hundreds of nm2 with mT at the AF domain on the right. Additionally, there is a zero-field line between domains. Since the NV-axis is essentially perpendicular to the sample surface, this zero-field line indicates that there is no field projection along the NV-axis. Therefore, it can be interpreted as the position at which the OOP component of the field is zero, which could define the position of the DW inside the FM stripe. 51, 48, 34. FM contours with mT are observed along one side of the zero-field line, while domains with opposite orientation of the field components are periodically separated by the zero-field line as shown in the image in Figure 3(f) for FM domains with mT.
@@ 2.4 Quantitative vector-field maps of a 3D spin texture
Furthermore, we record ODMR maps with a (111)-NV probe as shown in Figure 4(a). The field component projected along the NV-axis is spatially resolved over the sample surface. Given this crystallographic orientation of the tip, we are probing the = - component of the stray magnetic field. 2D maps of the in-plane (IP) stray field components can be reconstructed from the experimental map in Figure 4(a) by using the reverse propagation method in Fourier space 32, 8 (see section 9 in SI). The obtained and stray field maps are depicted in Figure 4(b) and (c), respectively. The maximum value for the static stray field is 8.0 mT at the boundary of the AF domains. The intensity of the stray field progressively decreases when moving far away from the boundary by hundreds of nm. Both IP and OOP components of the stray field are in the range from hundreds of T to 1mT on the AF domains. Even if we expect in a multilayered SAF structure with PMA a full compensation of the magnetization on the AF domains, there will be a small field sensed by the NV above the sample surface due to the different distances between the NV center and each of the FM blocks 12, 27. Additionally, from the images we observe a granular contrast, indicating magnetization variations (see section 10 in SI for a close-up ODMR map on the AF domains). These inhomogeneities or granularity are expected in sputtered multilayered samples with non-uniform thin film thicknesses or atomically flat surface that can induce local variation and non-compensation of the magnetization.
@@
Profiles taken across the region between the AF domains are plotted in Figure 4(g). For simplicity, we define the DW as the region where the OOP field component () is zero 48, 34, 51 (centred at distance=0 in the profiles). Since it is the region where there is a sign reversal of the stray field (as shown in the dual iso-B results). Note that it strictly corresponds to the actual DW center line (, -magnetization) only for DWs between two similar domains (AF-AF or FM-FM). Between a FM and an AF domain, however, we only consider the DW on the top FM block. Due to the extended FM stray field, the line is shifted towards the AF domain compared to the line. At 75 nm height above the surface, this shift amounts to several tens of nanometers according to our stray field simulation. From the profiles we can also see that above the DW, the IP field components and have a local minimum and a maximum, respectively. In contrast to what has been observed in ultra-thin film 48, 34, the profiles of the stray field intensity are not symmetric around the DW. We attribute this to the fact that we are imaging the DW at the top FM block with a collective contribution of the full multilayered structure to the total stray field, in agreement with reported MFM profiles.20, 21 The DW-type can in principle be inferred by fitting the field profiles using an analytical expression of the stray field above the DW under the approximation of a thin sample 48, i.e. a sample with magnetic thickness smaller than the sensor-sample distance. This approximation is not valid for the multilayered SAF investigated here since its magnetic thickness is comparable to the NV-sample distance. Previous micromagnetic simulations of multilayered structures demonstrated that FM domains are separated by Néel-type DWs at the top and bottom FM blocks with Bloch-type DWs in the middle of the structure 45. Based on this, we could expect a similar DW-type configuration; however, a direct determination of the DW type would require additional high-resolution studies, such as iteratively refined micromagnetic simulations or complementary local magnetometry techniques (for instance, Lorentz transmission electron microscopy), which are beyond the scope of the present work.
@@
Our vector-field maps are in good agreement with micromagnetic simulations. Figures 4(d)-(f) show images obtained from micromagnetic simulation of the stray field components (see methods and section 11 in SI) at a distance of 75 nm from the sample surface, in the range of the estimated . The sensitivity of the extracted DW position to the experimental uncertainty in can be evaluated by comparing the simulated stray field maps at heights of 36 nm and 75 nm. We find that varying the NV-sample distance within this range results in a change of approximately 3 nm in the absolute DW position, corresponding to a relative sensitivity of 8% with respect to the total variation in . This value depends on the specific position of the profile used to extract the DW position; nevertheless, the variation is expected to remain of the same order even with a more rigorous statistical analysis. From the simulations we observed stray magnetic field of approximately 12 mT. The absolute values of the stray field are higher than the experimental values, suggesting that the magnetic parameters used to initially define the simulated structure need to be optimized as well as the determination of the NV-sample distance. However, the simulated and experimental stray-field maps are qualitatively in good agreement with the stray field values in the same order of magnitude, meaning that the remanent relaxed state achieved in the simulation can describe the magnetic texture of the multilayered SAF. Particularly, the granularity inducing a non-compensated magnetization on the AF domains is also reproduced in the simulated images. The FM domains along the stripe between the two AF domains are also well reproduced by the micromagnetic simulation. To visualize the domain structure in a 2D image, Figure 4(h) displays a transparent overlap of the simulated OOP magnetization of the top and second FM blocks (see section 11 in SI). The FM domains around the DWs are formed due to an alternating left and right lateral shift between DWs in adjacent FM blocks, leading to a vertical alignment of the magnetic moments along the stacks and giving rise to a 3D spin texture as illustrated in the inset. This lateral shift has been previously reported as driven by the dipolar interaction against the AF-exchange interaction between the FM blocks with a periodic reversal of the FM cores to minimize the magnetostatic energy produced around the DWs, and with a constant DW-shift of nm along the FM stripe between AF domains for a multilayered structure with and . 20 In another study, Kiselev et al. 26 describe, within their micromagnetic model, that the DW-shift configuration cannot be stabilized only by the interplay between the magnetostatic and the AF-exchange interaction. The stability depends also on DW pinning (grain boundaries, defects or inhomogeneities) that give rise to nucleation sites for one-dimensional FM stripes with fixed width (DW-shift). Remarkably, in the ODMR vector field maps, the periodic modulation of the stray field as a consequence of the DW-shift that reduce the magnetostatic energy is quantitatively and spatially resolved in both the experimental and the simulated IP field maps in Figure 4 (see black arrows). Additionally, the observed granularity or inhomogeneity can favor the stabilization of DW shifts through the DW pinning. We found that the lateral DW-shift is non-uniform along the FM stripes, it varies in the range nm (see section 11 in SI), generating FM domain cores with different nanometer sizes and wave-shaped DWs.
@@ 2.5 Sensing different components of the magnetic noise
Finally, in order to understand the spin dynamics of the magnetic texture in the multilayered SAF, we performed relaxometry measurements to detect spin noise. is very sensitive to magnetic fluctuations perpendicular to the NV-axis and with frequencies near the NV resonance frequency 7. To investigate spin noise, we performed all-optical measurements 31 using the pulse sequences in Figure 5(a). We use a green laser pulse for initialization of the NV-spin states in , wait for some time without laser illumination, and then a second green laser pulse is sent to read out the final spin state. Figure 5(b) shows the measurements on the (111)-oriented diamond tip. We first measure before approaching the sample surface, obtaining s. This value is expected for diamond tips with shallow NV (nominal implantation depth of 10 nm), where the relaxation time is mostly affected by paramagnetic impurities on the diamond surface 41. After the sample approach and positioning the tip on the AF domain and then on the FM stripe, we observe a drop in by almost one and two orders of magnitude, respectively. We obtain s when the tip is in the AF domain and even shorter s on the FM stripe. Recent calculations of the spin-wave spectrum in multilayered SAFs 24 predict GHz optical and acoustic spin-wave modes, with the dispersion relation and number of modes dependent on the thickness and number of FM blocks. Additionally, our results are in agreement with the results reported by Finco et al 12 on the study of magnetic noise on AF DWs. Thus, we attribute these experimental observations to the presence of magnetic noise in the GHz range, most likely due to thermal magnons with frequencies near the NV spin transition frequency. is shorter on the FM stripes than on the AF domains, probably due to the confinement of thermal magnonic modes with a stronger noise or frequency that matches the electron spin resonance frequency of the NV at 2.87 GHz.
@@
Figure 5(c) shows a comparison of on NV diamond scanning probes with different crystallographic orientations (see section 12 in SI). The change in is essentially the same for the three diamond probes, indicating the existence of magnetic noise with components perpendicular to the respective NV-axis. These measurements suggest that the PL quenching observed on the FM stripes between AF domains during the acquisition of NV-PL maps could have contributions from both magnetic noise and spin-mixing. For the measurement with the (100)-NV probe, the dominant contribution to the NV-PL quenching comes probably from spin-mixing, as demonstrated by local ODMR measurements. A similar effect is expected for the (110)-NV tip. For the (111)-NV tip, the IP components resolved in the ODMR maps could in principle produce spin mixing. However, there is no evidence of spin-mixing from local ODMR measurements. Hence, the magnetic noise is the dominant contribution that produces quenching in the NV-PL images. Under this condition, it should be possible to observe an improvement in the PL contrast in the NV-PL images when the laser power is decreased below saturation, as observed on samples with AF DW 12. However, NV-PL quenching images recorded with different laser powers (section 13 in SI) show no improvement in PL contrast, indicating that the field of the FM core along the DW could potentially play a role in the relation between the spin-polarization rate given by the laser power and the spin relaxation rate defined by .
In the model reported by Finco et al 12, the contrast in NV-PL quenching maps is described as a competition between the spin relaxation rate and the optical pumping rate, without considering the effect of the sample stray field. In this case, states of the NV-probe are degenerate. In our multilayered SAF, in addition to the GHz magnetic noise, the FM cores generate a local magnetic field that varies spatially along the boundaries between AF domains. As a result, the states are no longer degenerate and therefore have different occupation probabilities. This likely alters the initial conditions and symmetry assumed in the two-level model, which leads to different coupling conditions between the NV spin transitions and the magnetic noise. Consequently, the improvement of the PL contrast at reduced laser power is not observed in our mixed-order structure.
@@ 3 Conclusion
Quantitative vector-field maps combined with micromagnetic simulations allow us to explore interlayer coupling effects in the multilayered structure through direct magnetic imaging of DW shifts and the periodic modulation of the IP stray field components produced by the formation of nanometric FM cores around undulated DWs. These results provide validation on an experimental and micromagnetic simulation level and further insights into existing models about the observed 3D spin texture. Additionally, local relaxometry measurements reveal the presence of GHz-range magnetic noise on both the AF domains and the DWs with FM cores.
@@ 4.1 Sample preparation and characterization
The multilayered SAF was prepared at room temperature by dc magnetron sputtering deposition (ATC 2200 series sputter tool from AJA International, Inc.) with a base pressure of Pa, using Ar sputter gas with 0.4 Pa pressure. Silicon substrates with a thermal 100 nm SiO2 layer were used, on which a 1.5 nm thick Ta adhesion layer10, 20 nm Pt seed layer for promoting (111) texture59, and 3 nm Pt cap against oxidation were deposited.
@@ 4.2 NV-SPM measurements
Qualitative measurements (NV-PL maps and ) were mostly performed in a new home-built setup, which combines tuning-fork AFM (attocube tip and sample positioner and scanner) with confocal optical microscopy. NV centers in the diamond probes are optically pumped with a 515 nm laser (Toptica, iBEAM-SMART-515-S) and readout by detecting the NV-PL signal after passing through a long-pass 650 nm filter (Thorlabs, FELH0650). The signal is collected by using dual single photon counting modules (Excelitas, SPCM-AQRH-14) mounted in a Hanbury Brown and Twiss geometry for photon autocorrelation measurements. The excitation power of the green laser is 130-300 W, measured after the objective lens (Olympus, MPLFLN100x) that focus (collect) the light on (from) the tip. We use Qzabre quantum scanning tips with sensitivity Q6-Q7 and nominal NV implantation depth of 10 nm. Quantitative measurements (ODMR and dual-iso B) were carried out in a commercial Quantum Scanning NV Microscope (QSM, Qzabre AG) equipped with an electromagnet to apply magnetic field in arbitrary directions. All the NV-SPM images were acquired using the non-contact amplitude modulation mode. The NV-sample distance, as well as, the orientation of the NV-axis in the laboratory frame are determined using the method described in reference 23, on a 1 m-wide Ta/Pt/Co(0.6nm)/Pt wire with perpendicular magnetic anisotropy as a calibration sample (see SI for more details).